# Contemporaneous and lagged spillovers across crude oil, carbon emission allowance, climate change, and agriculture futures markets: Evidence from the $R^{2}$ decomposed connectedness approach

### v1

@@ Abstract
In this paper, we examine the dynamic spillovers among the crude oil, carbon emission allowance, climate change, and agricultural markets. Adopting a novel $R^{2}$ decomposed connectedness approach, our empirical analysis reveals several key findings. The overall TCI dynamics have been mainly dominated by contemporaneous dynamics rather than the lagged dynamics. We also find climate change has significant spillovers to other markets. Moreover, there are heterogeneous spillover effects among agricultural markets. Specially, corn is the biggest risk contributor to this system, while barley is the major risk receiver of shocks.
@@
Dynamic connectedness , Crude oil , Carbon emission allowance , Climate change , Agricultural commodities , $R^{2}$ decomposition
@@ 1 Introduction
As global warming intensifies and extreme weather becomes more frequent, concerns about potential economic losses and impacts on social development have increased, leading to widespread attention to climate risks (Venturini, 2022; Kemp et al., 2022; Pham and Kamal, 2024; Khalfaoui et al., 2024). Severe physical climate risks affect the cultivation of agricultural products. Additionally, carbon dioxide also exerts a negative impact on agricultural and food markets (Khalfaoui et al., 2024). Climate change influences the extraction and transportation of fossil energy. Volumetric oil production declines with significant growth of greenhouse gas (Masnadi and Brandt, 2017). However, energy production and agricultural activities can lead to environmental pollution, worsening climate conditions and intensifying global warming (Jin et al., 2023).
@@
Since agricultural and energy markets have always been fundamentally connected (Du et al., 2011; Shao et al., 2024), climate change has added new complexity to this interconnection. Moreover, the financialization of commodities has enriched investment instruments but has also heightened the risk of contagion. Owing to the strong interconnections within the financial system, a shock in one sector can affect other sectors. Furthermore, the contagion effect can extend to the entire financial market. Hence, the research of spillover effect and information transmission among climate change and commodity markets continues to expand, which is crucial for assessing contagion risks, price forecasting, and portfolio management. Jin et al. (2023) report the dynamic spillover among geopolitical risk, climate risk, and energy markets from an international perspective. Lin et al. (2024) observe simultaneous opposite correlations between climate change attention and crude oil markets.
@@
In response to the crucial climate issues, a carbon emission allowance trading market has been developed. The market provides economic incentives for companies to adopt cleaner technologies and energy sources to reduce emission costs, helping to mitigate climate change. It should be noted that fossil energy and agricultural markets have important impacts on the carbon market, which will influence carbon demand, supply, and prices (Gong et al., 2021). Chen et al. (2024) provide empirical evidence that the risk is transferred from the international gas and crude oil markets to the Chinese carbon market. Younis et al. (2024) investigates contagion risk among climate change, renewable energy, technological innovation, and G-17 banking stock markets. They find significant risk spillover and connectedness.
Recently, the interdependency among energy and agricultural commodity markets has been widely concerned (Wei et al., 2023). However, existing research has mainly focused on overall spillover effects as well as long-term and short-term frequency spillovers. There is relatively limited empirical analysis that decomposes the spillover effects into contemporaneous and lagged periods (Balli et al., 2023). Furthermore, there is a lack of detailed exploration of the risk spillover among agricultural markets in existing studies. Hence, we examine the dynamic spillover among carbon emission allowance, climate change, crude oil, and 16 agricultural futures.
The contribution of our study is as follows. First, we apply a novel $R^{2}$ decomposed connectedness approach proposed by Balli et al. (2023), which could distinguish the contemporaneous and lagged spillovers. Studying contemporaneous and lagged spillover effects in financial markets is crucial for investment decisions, risk management, and ensuring market stability and healthy development. Second, we conduct detailed analysis on how different agricultural products respond to climate change, carbon emission allowances, and crude oil fluctuations, thereby offering targeted risk management strategies. Third, this paper expands the study on factors influencing commodity prices. It reveals dynamic correlations among crude oil, carbon emission allowance, climate change, and agricultural markets.
@@
The remainder of this paper is organized as follows. Section 2 outlines the methodological framework. Section 3 describes the data. Section 4 presents the empirical results and robustness checks. Section 5 concludes the study.
@@ 2 Methodology
Building upon a novel approach for assessing $R^{2}$ decomposed connectedness (Balli et al., 2023; Naeem et al., 2024), this study identifies both contemporaneous and lagged idiosyncratic spillover effects within the crude oil, carbon emission allowance, climate change, and agricultural futures markets. This addresses the earlier research that primarily focused on contemporaneous spillovers, marking a significant expansion of the analysis. As detailed by Balli et al. (2023), the $R^{2}$ decomposed connectedness framework utilizes a VAR($p$) model to incorporate contemporaneous effects:
@@
Typically, the cumulative $R^{2}$ contributions from bivariate linear regressions (BLRs) correspond with the multivariate linear regression (MLR)’s goodness-of-fit only when the variables on the right-hand side (RHS) are not correlated with one another. It is crucial, therefore, to establish a transformation that converts the correlated series $\bm{x}_{k,t}$ [^1] The vector $\bm{x}_{k,t}$ corresponds to $\bm{x}_{t}$but omits the variable on the left-hand side. into a set of orthogonal series. This transformation is facilitated by implementing principal component analysis, where the number of latent factors matches the quantity of RHS variables. As a result, the $R^{2}$ decomposition of an MLR can be formulated as follows:
@@
Specifically, the $TO_{i}$ index quantifies the proportion of variance in all LHS variables that is attributed to variable $i$. In contrast, the $FROM_{i}$ index measures how much the RHS variables collectively account for the variance in variable $i$. From the perspective of spillover effects, $NET_{i}$ index indicates whether a variable is a net transmitter or receiver of shocks. If series $i$ has a positive (negative) $NET$ value, it is deemed a net transmitter (receiver) of shocks, suggesting it influences other variables more (less) than it is influenced by them. Analogously, the net pairwise directional connectedness ($NPDC_{ij}$), which works in the same manner as the $NET$ connectedness but operates on a mutual, two-way level between pairs of variables. A positive (negative) $NPDC_{ij}$ indicates that variable $j$ accounts for more (less) of the variation in variable $i$ than vice versa.
Table 1: Summary statistics
OilB CarE CliC Soyb Corn Wheat Rice Sugar Cotto Coffe Cocoa Oats Soym Soyo Cattl Milk Lhog Juice Barl
Mean 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.001 0.000 0.001 0.001 0.000 0.000 0.000 0.001 0.000 0.000 0.002 0.000
Variance 0.001 0.001 0.000 0.000 0.001 0.001 0.000 0.000 0.001 0.000 0.001 0.001 0.005 0.001 0.000 0.000 0.001 0.001 0.000
Skewness -0.730*** -0.577*** -0.020 -1.095*** -1.898*** -0.358*** -1.354*** -0.281*** -3.375*** 0.433*** -1.083*** -2.631*** -0.459*** -0.653*** 0.283*** 3.442*** -2.205*** -0.762*** 0.273***
Ex.Kurtosis 3.370*** 5.491*** 1.779*** 7.565*** 19.347*** 11.338*** 15.520*** 1.780*** 46.721*** 2.946*** 71.817*** 31.460*** 13.158*** 2.701*** 11.247*** 68.055*** 32.473*** 3.566*** 4.315***
JB 433.988*** 1012.642*** 101.843*** 1994.993*** 12503.623*** 4151.900*** 7983.829*** 112.122*** 71680.628*** 303.292*** 166056.293*** 32726.651*** 5596.583*** 289.566*** 4079.262*** 150503.046*** 34545.979*** 483.635*** 608.492***
ERS -10.328*** -8.316*** -9.559*** -12.933*** -12.054*** -11.440*** -12.706*** -5.708*** -13.173*** -7.954*** -1.935* -13.461*** -13.324*** -$1**$1.856*** -8.076*** -11.575 *** -5.442 *** -9.041***
Kendall’s $\rho$ OilB CarE CliC Soyb Corn Wheat Rice Sugar Cotto Coffe Cocoa Oats Soym Soyo Cattl Milk Lhog Juice Barl
OilB 1.000*** 0.065*** 0.099*** 0.172*** 0.125*** 0.119*** 0.089*** 0.136*** 0.179*** 0.041 0.071*** 0.076*** 0.029 0.205*** 0.065*** 0.011 0.041 -0.008 0.014
CarE 0.065*** 1.000*** 0.083*** 0.034 -0.044 -0.035 -0.007 0.047 0.030 0.020 0.028 0.008 0.005 0.073*** 0.034 -0.007 0.013 -0.011 -0.044
CliC 0.099*** 0.083*** 1.000*** 0.039 -0.006 -0.004 -0.003 0.044 0.110*** 0.036 0.051** 0.043 -0.008 0.065*** 0.072*** -0.020 0.028 0.022 -0.036
Soyb 0.172*** 0.034 0.039 1.000*** 0.347*** 0.265*** 0.122*** 0.136*** 0.094*** 0.102*** 0.003 0.161*** 0.379*** 0.351*** 0.055** 0.031 0.036 0.046 0.018
Corn 0.125*** -0.044 -0.006 0.347*** 1.000*** 0.385*** 0.112*** 0.136*** 0.090*** 0.094*** -0.024 0.201*** 0.159*** 0.190*** 0.034 0.012 0.041 0.004 0.036
Wheat 0.119*** -0.035 -0.004 0.265*** 0.385*** 1.000*** 0.131*** 0.148*** 0.125*** 0.139*** 0.011 0.204*** 0.135*** 0.179*** 0.001 0.024 0.048** 0.019 0.035
Rice 0.089*** -0.007 -0.003 0.122*** 0.112*** 0.131*** 1.000*** 0.060** 0.047 0.056** 0.020 0.063** 0.018 0.104*** 0.015 0.019 -0.016 0.008 0.015
Sugar 0.136*** 0.047 0.044 0.136*** 0.136*** 0.148*** 0.060** 1.000*** 0.150*** 0.158*** 0.066*** 0.132*** 0.066*** 0.120*** 0.005 0.030 0.028 0.047 0.027
Cotto 0.179*** 0.030 0.110*** 0.094*** 0.090*** 0.125*** 0.047 0.150*** 1.000*** 0.124*** 0.042 0.109*** 0.071*** 0.104*** 0.057** -0.034 0.059** 0.041 -0.020
Coffe 0.041 0.020 0.036 0.102*** 0.094*** 0.139*** 0.056** 0.158*** 0.124*** 1.000*** 0.089*** 0.107*** 0.060** 0.077*** 0.016 0.004 0.033 0.022 0.036
Cocoa 0.071*** 0.028 0.051** 0.003 -0.024 0.011 0.020 0.066*** 0.042 0.089*** 1.000*** 0.024 -0.001 0.022 0.067*** -0.051** -0.005 -0.047 0.010
Oats 0.076*** 0.008 0.043 0.161*** 0.201*** 0.204*** 0.063** 0.132*** 0.109*** 0.107*** 0.024 1.000*** 0.099*** 0.122*** 0.039 -0.041 0.025 0.011 -0.002
Soym 0.029 0.005 -0.008 0.379*** 0.159*** 0.135*** 0.018 0.066*** 0.071*** $0*$001 0.099*** 1.000*** 0.021 0.021 -0.003 0.004 -0.012 -0.027
Soyo 0.205*** 0.073*** 0.065*** 0.351*** 0.190*** 0.179*** 0.104*** 0.120*** 0.104*** 0.077*** 0.022 0.122*** 0.021 1.000*** 0.066*** 0.015 0.059** 0.043 0.060**
Cattl 0.065*** 0.034 0.072*** 0.055** 0.034 0.001 0.015 0.005 0.057** 0.016 0.067*** 0.039 0.021 0.066*** 1.000*** 0.039 0.064*** 0.006 0.010
Milk 0.011 -0.007 -0.020 0.031 0.012 0.024 0.019 0.030 -0.034 0.004 -0.051** -0.041 -0.003 0.015 0.039 1.000*** 0.016 0.017 -0.007
Lhog 0.041 0.013 0.028 0.036 0.041 0.048** -0.016 0.028 0.059** 0.033 -0.005 0.025 0.004 0.059** 0.064*** 0.016 1.$0**$010 -0.061**
Juice -0.008 -0.011 0.022 0.046 0.004 0.019 0.008 0.047 0.041 0.022 -0.047 0.011 -0.012 0.043 0.006 0.017 -0.010 1.000*** 0.004
Barl 0.014 -0.044 -0.036 0.018 0.036 0.035 0.015 0.027 -0.020 0.036 0.010 -0.002 -0.027 0.060** 0.010 -0.007 -0.061** 0.004 1.000***
Notes: JB and ERS show the results of normality test and unit root test, respectively. * indicates 10% significance level, ** indicates 5% significance level, *** indicates 1% significance level.
@@ 3 Data
We select a dataset encompassing various futures markets, including crude oil, carbon emission allowance, climate change, and agriculture, totaling 19 different futures contracts.
The data spans from December 15, 2020, to November 30, 2023, and consists of daily closing prices of continuous contracts, comprising 773 data observation points in total. To ensure data consistency, missing data points are filled with the preceding day’s data. The data is sourced from Wind and Bloomberg.
The starting date of December 15, 2020, was specifically chosen as it marks the launch of the climate change futures.
@@
In the oil category, we included the Brent Oil Future, which is actively traded on the ICE Exchange. For carbon emission allowance, our dataset features the EUA Carbon traded on ICE Futures Europe. In the climate segment, we have incorporated the MSCI World Climate Change Futures, listed on the ICE Exchange, marking an essential inclusion considering the global focus on climate-related financial products.
@@
The agricultural futures in our dataset are quite extensive. They include Soybean, Soybean meal, Soybean oil, Corn, Wheat, Oats, and Rough Rice Futures, all traded on the CBOT Exchange. The ICE Exchange is represented with Sugar #11 Futures, Cotton #2 Futures, Orange juice, Cocoa, and Coffee C Futures; notably, the Coffee C futures contract on ICE serves as the global benchmark for Arabica coffee. Additionally, Live cattle, Milk, Lean hog futures from the CME Exchange are part of our selection. Reflecting the significance of Australia in the global agricultural market, particularly in barley, we also consider Barley Futures listed on the Australian Securities Exchange.
@@
For the sake of simplicity in the presentation of tables and figures, we employ the abbreviations OilB, CarE, CliC, Soyb, Rice, Sugar, Cotto, Coffe, Soym, Soyo, Cattl, Lhog, Juice, and Barl to denote the futures Brent Oil, EUA Carbon, MSCI World Climate Change, Soybean, Rough Rice, Sugar #11, Cotton #2, Coffee, Soybean Meal, Soybean Oil, Live Cattle, Lean Hog, Orange Juice, and Barley, respectively.
@@
Table. 1 reports the descriptive statistic. We find that except coffee, cattle, milk, and barley, other variables are left skewed. Besides that, with the exception of climate change, sugar and soybean oil, others are leptokurtic. The JB and ERS tests indicate that all series are non-normally distributed and significantly stationary. Moreover, the Kendall rank correlation coefficients suggest that most variables are positively correlated.

### v2

# Contemporaneous and lagged spillovers between agriculture, crude oil, carbon emission allowance, and climate change

@@ Abstract
In this paper, we examine the contemporaneous and lagged spillovers among the agricultural, crude oil, carbon emission allowance, and climate change markets. Adopting the $R^{2}$ decomposed connectedness approach, our empirical analysis reveals several key findings. First, the overall total connectedness index (TCI) dynamics have been mainly dominated by contemporaneous effects. Second, there are heterogeneous spillover effects among agricultural markets. Specially, corn, soybean meal, and wheat are the major risk transmitters to this system, while barley, cocoa, and lean hog are the main risk receivers of shocks. Third, we also find that climate change has significant spillovers to other markets.
@@
Dynamic connectedness , Agricultural commodities, Crude oil , Carbon emission allowance , Climate change , $R^{2}$ decomposition
@@ 1 Introduction
The agricultural market is crucial to the national economy, which is susceptible to external shocks. Intensifying global warming and extreme weather have heightened attention to climate risks (Venturini, 2022; Kemp et al., 2022; Pham and Kamal, 2024; Khalfaoui et al., 2024), severely impacting crop cultivation. Additionally, carbon dioxide also exerts a negative impact on agricultural markets (Khalfaoui et al., 2024). Agricultural and energy markets have long been connected (Du et al., 2011), with climate change adding complexity by affecting fossil energy extraction and transport. Volumetric oil production declines with significant growth of greenhouse gas (Masnadi and Brandt, 2017). However, energy production and agriculture can worsen pollution, climate conditions, and global warming (Jin et al., 2023).
@@
Owing to the strong interconnections within the financial system, a shock in one sector can affect other sectors (Wang et al., 2020, 2022). Furthermore, the contagion effect can spread across the financial market, making spillover research among agricultural, energy, and other markets vital for assessing risks, price forecasting, and portfolio management (Wei et al., 2023). Jin et al. (2023) report the dynamic spillover among geopolitical risk, climate risk, and energy markets from an international perspective. Lin et al. (2024) observe simultaneous opposite correlations between climate change attention and crude oil markets.
@@
In response to the crucial climate issues, the carbon emission allowance trading market has been developed, helping to mitigate climate change. It should be noted that fossil energy and agricultural markets have important impacts on the carbon market, which will influence carbon demand, supply, and prices (Gong et al., 2021). Chen et al. (2024) find that the risk is transferred from the international gas and crude oil markets to the Chinese carbon market. Younis et al. (2024) invesitgate contagion risk among climate change, renewable energy, technological innovation, and G-17 banking stock markets. They find significant risk spillover and connectedness.
There are various measures to investigate risk transmission in agricultural and other commodity markets, including DCC-GARCH (Luo and Ji, 2018), CoVaR (Kumar et al., 2021), TVP-VAR-DY (Zhu et al., 2024), quantile VAR model (Tiwari et al., 2022; Wei et al., 2023). Recently, Balli et al. (2023) and Naeem et al. (2024) introduce the $R^{2}$ decomposed connectedness approach, efficiently splitting connectedness into contemporaneous and lagged components while ensuring data standardization. This method has since been applied to risk spillover analysis in financial markets (Li et al., 2024; Liu et al., 2024).
Commodity financialization boosts liquidity and market efficiency, but also raises contagion risk (Goldstein and Yang, 2022; Bianchi et al., 2020). Hence, we examine the spillover among agriculture, crude oil, carbon emission allowance, and climate change futures markets via the $R^{2}$ decomposed connectedness approach. Our study contributes the following: First, we apply the $R^{2}$ decomposed connectedness to accurately identify the immediate and delayed information transmissions between commodity markets, providing a comprehensive understanding of how risks are transmitted across different time dimensions. Second, this paper deepens the understanding of risk transmission between commodity markets by expanding the study on factors influencing commodity prices. Third, we conduct detailed analysis on how different agricultural products respond to other futures fluctuations, offering targeted risk management strategies for investors and policymakers.
@@
The paper is organized as follows: Section 2 outlines the methodology, Section 3 describes the data, Section 4 presents empirical results and robustness checks, and Section 5 concludes.
@@ 2 Methodology
Building on the $R^{2}$ decomposed connectedness approach (Balli et al., 2023; Naeem et al., 2024), this study identifies both contemporaneous and lagged idiosyncratic spillovers in agricultural and other futures markets, expanding earlier research that focused solely on contemporaneous effects. As detailed by Balli et al. (2023), the $R^{2}$ decomposed connectedness framework utilizes a VAR($p$) model to incorporate contemporaneous effects:
@@
Typically, the cumulative $R^{2}$ contributions from bivariate linear regressions (BLRs) correspond with the multivariate linear regression (MLR)’s goodness-of-fit only when the variables on the right-hand side (RHS) are not correlated with one another. It is crucial to establish a transformation that converts the correlated series $\bm{x}_{k,t}$ [^1] The vector $\bm{x}_{k,t}$ corresponds to $\bm{x}_{t}$but omits the variable on the left-hand side. into a set of orthogonal series. This transformation is facilitated by implementing principal component analysis, where the number of latent factors matches the quantity of RHS variables. As a result, the $R^{2}$ decomposition of an MLR can be formulated as follows:
@@
The $TO_{i}$ index quantifies the variance in LHS variables due to $i$, while the $FROM_{i}$ index measures how much RHS variables explain the variance in $i$. The $NET_{i}$ index shows whether $i$ is a net transmitter or receiver of shocks. If series $i$ has a positive (negative) $NET$ value, it is deemed a net transmitter (receiver) of shocks, suggesting it influences other variables more (less) than it is influenced by them. Similarly, net pairwise directional connectedness ($NPDC_{ij}$) operates like $NET$ but between pairs. A positive (negative) $NPDC_{ij}$ means $j$ explains more (less) of $i$’s variation than vice versa.
@@ 3 Data
Our dataset includes 19 futures contracts in agriculture, crude oil, carbon emission allowances, and climate change, spanning from December 15, 2020, to November 30, 2023, with 773 daily closing price observations. The starting date, December 15, 2020, marks the launch of climate change futures. Missing data points are filled with the previous day’s data for consistency. The data is sourced from Wind and Bloomberg. We compute logarithmic returns of futures prices as $\ln\left(\frac{p_{i,t}}{p_{i,t-1}}\right)$, where $p_{i,t}$ is the daily closing price of the $i$-th future at time $t$, as shown in Fig. 1.
@@
We include Brent Oil Futures from the ICE Exchange in the oil category. For carbon emission allowance, our dataset features the EUA Carbon traded on ICE Futures Europe. In the climate segment, we include the MSCI World Climate Change Futures from the ICE Exchange, a key addition given the global focus on climate-related financial products.
@@
Our dataset includes a wide range of agricultural futures: Soybean, Soybean meal, Soybean oil, Corn, Wheat, Oats, and Rough Rice, all traded on the CBOT Exchange. The ICE Exchange is represented with Sugar #11 Futures, Cotton #2 Futures, Orange juice, Cocoa, and Coffee C Futures; notably, the Coffee C futures contract on ICE serves as the global benchmark for Arabica coffee. Additionally, Live cattle, Milk, and Lean hog futures from the CME Exchange are part of our selection. Given Australia’s significance in the global barley market, we include Barley Futures from the Australian Securities Exchange.
Figure 1: Returns.
Table 1: Summary statistics.
Mean Variance Skewness Ex.Kurtosis JB ERS ADF PP
OilB 0.001 0.001 -0.730*** 3.370*** 433.988*** -10.328*** -10.645*** -678.469***
CarE 0.001 0.001 -0.577*** 5.491*** 1012.642*** -8.316*** -9.636*** -817.237***
CliC 0.000 0.000 -0.020 1.779*** 101.843*** -9.559*** -8.786*** -683.829***
Soyb 0.000 $0$1.095*** 7.565*** 1994.993*** -12.933*** -8.913*** -723.122***
Corn 0.000 0.001 -1.898*** 19.347*** 12503.623*** -12.054*** -8.895*** -810.834***
Wheat 0.000 0.001 -0.358*** 11.338*** 4151.900*** -11.440*** -8.625*** -768.58***
Rice 0.000 $0$1.354*** 15.520*** 7983.829*** -12.706*** -10.135*** -625.387***
Sugar 0.001 $0$0.281*** 1.780*** 112.122*** -5.708*** -10.479*** -760.936***
Cotto 0.000 0.001 -3.375*** 46.721*** 71680.628*** -13.173*** -8.824*** -720.521***
Coffe 0.001 0.000 0.433*** 2.946*** 303.292*** -7.954*** -8.99*** -761.17***
Cocoa 0.001 0.001 -1.083*** 71.817*** 166056.293*** -1.935* -11.786*** -801.937***
Oats 0.000 0.001 -2.631*** 31.460*** 32726.651*** -13.461*** -9.374*** -665.873***
Soym 0.000 0.005 -0.459*** 13.158*** 5596.583*** -13.324*** -12.271*** -917.347***
Soyo 0.000 0.001 -0.653*** 2.701*** 289.566*** -11.401*** -7.842*** -672.379***
Cattl 0.001 0.000 0.283*** 11.247*** 4079.262*** -11.856*** -8.694*** -767.425***
Milk 0.000 0.000 3.442*** 68.055*** 150503.046*** -8.076*** -8.876*** -779.297***
Lhog 0.000 0.001 -2.205*** 32.473*** 34545.979*** -11.575*** -9.838*** -754.072***
Juice 0.002 0.001 -0.762*** 3.566*** 483.635*** -5.442*** -11.262*** -678.677***
Barl 0.000 0.000 0.273*** 4.315*** 608.492*** -9.041*** -7.717*** -884.229***
  • 1.
    Notes: JB represents the Jarque-Bera normality test statistic; ERS, ADF, and PP show the results of the unit root tests. * indicates 10% significance level, ** indicates 5% significance level, *** indicates 1% significance level.
@@
To simplify the presentation of tables and figures, we employ the abbreviations OilB, CarE, CliC, Soyb, Rice, Sugar, Cotto, Coffe, Soym, Soyo, Cattl, Lhog, Juice, and Barl to denote the futures Brent Oil, EUA Carbon, MSCI World Climate Change, Soybean, Rough Rice, Sugar #11, Cotton #2, Coffee, Soybean Meal, Soybean Oil, Live Cattle, Lean Hog, Orange Juice, and Barley, respectively.
Figure 2: Kendall correlations. Notes: * Significant at 10%; ** significant at 5%; *** significant at 1%.
@@
Table 1 reports the descriptive statistics. Except coffee, cattle, milk, and barley, other variables are left skewed. Besides that, with the exception of climate change, sugar, and soybean oil, others are leptokurtic.The JB, ERS, ADF, and PP tests indicate that all series are non-normally distributed and stationary. Moreover, the Kendall rank correlation coefficients presented in Fig. 2 suggest that most variables are positively correlated.